OpenStudy (anonymous):

WILL GIVE MEDAL!!!! If tan(theta) = 3/4 and theta is an acute angle, find cos(theta)

OpenStudy (anonymous):

HInt;- \[\sec^2(\theta) - \tan^2(\theta) = 1\] from above determine sec(theta)

OpenStudy (anonymous):

remember \[\cos(\theta) = \frac{ 1 }{ \sec(\theta) }\]

OpenStudy (anonymous):

I'm just lost. I don't even know what answer I am supposed to have

OpenStudy (anonymous):

in ques value of tan(theta) is given from this u can determine sec(theta)

OpenStudy (anonymous):

hmmm how would I do that? find sin?

OpenStudy (anonymous):

@niksva

OpenStudy (anonymous):

no tan(theta) = 3/4 is given sec^2(theta)-tan^2(theta) = 1 now plug in the value of tan(theta) and see what will come?

OpenStudy (anonymous):

x=sqrt19/4?

OpenStudy (anonymous):

@niksva

OpenStudy (anonymous):

sec^2(theta) - (3/4)^2 = 1 sec^2(theta) - 9/16 = 1 sec^2(theta) = 1 +(9/16) sec^2(theta) = (16+9)/16 sec^2(theta) = 25/16 sec(theta) = 5/4 now u can determine cos(theta) ??????????

OpenStudy (anonymous):

cos(theta) = 4/5? So 4/5 is my answer

OpenStudy (anonymous):

@niksva

OpenStudy (anonymous):

yeah!!!

OpenStudy (anonymous):

What about: If sin(theta)=12/13 and theta is a acute angle, find cot(theta)

OpenStudy (anonymous):

i have shown all the steps of previous answer now u have to show me your work first

OpenStudy (anonymous):

I know that cot=1/tan and that tan=sin/cos. so tan=(12/13)/cos

OpenStudy (anonymous):

from sin(theta) , determine cosec(theta) first? then use cosec^2(theta) - cot^2(theta) = 1

OpenStudy (anonymous):

this is all new to me. I don't know how to do that

OpenStudy (anonymous):

@niksva why can't you just show me step by step?

OpenStudy (anonymous):

i have shown u the steps of the previous ques u need to try it out first because i m here to help only and this time show your steps systematically

OpenStudy (anonymous):

but I have no clue how to do it. you showed me steps to finding cos, not cot

OpenStudy (anonymous):

@niksva

OpenStudy (anonymous):

value of sin(theta) is given tell me what is cosec(theta) ?

OpenStudy (anonymous):

csc=13/12

OpenStudy (anonymous):

@pgpilot326 please help

OpenStudy (gorv):

\[\tan \theta=\frac{ 3 }{ 4 }=\frac{ perpendicular }{ base }\]

OpenStudy (gorv):

hypotenuse= root(perpendicular ^2 + base^2)=root(3^2+4^2)=root(9+16)=root25 =5 \[\cos \theta = \frac{ base }{ hypotenuse}=\frac{ 4 }{ 5}\]

OpenStudy (gorv):

@washcaps

OpenStudy (anonymous):

hey @gorv I actually answered the first question already as 4/5 is my answer. I need help on the second

OpenStudy (anonymous):

draw a pic, then it's easier to see